OpenStudy (anonymous):

What is the distance between the points (-4, -9) and (-8, -5)?

Parth (parthkohli):

First find the distance between the x coordinates, then find the distance between y coordinates.

Parth (parthkohli):

When you actually draw two lines from x1 to x2 and then y1 to y2, you'll notice that the hypotenuse is the distance :)

OpenStudy (shane_b):

\[d=\sqrt{(x_2-x_1)^2+(y_2-y_1)^2}\]

Parth (parthkohli):

No need to learn all those formulas haha they sometimes really get you confused :)

OpenStudy (shane_b):

That's an easy one though...no drawing involved :)

Parth (parthkohli):

Haha lol you must not draw.

OpenStudy (shane_b):

I draw like a 4 year old!

Parth (parthkohli):

Like the distance between -4 and -8 is 4. Distance between -9 and -4 is 4. The hypotenuse can be found by \(c^2 = 4^2 + 4^2\)

Parth (parthkohli):

\(c^2 = 32\) \( \color{Black}{\Rightarrow c = \sqrt{32} }\) Simplify it.

OpenStudy (anonymous):

you can make it become a vector then find the length of that vector \[\sqrt{(x2-x1)^2 + (y2-y1)^2}\] = distance

OpenStudy (shane_b):

We already covered that :)